### v3

@@ 5 Conclusion
The primary focus of our experimental investigations revolved around two fundamental medical imaging tasks i.e, classification and segmentation. Significantly, our research findings indicate that the shift from a transfer-learning-based solely on supervised methodology to a machine-supervised framework, which we named S4MI (Self-Supervision and Semi-Supervisio for Medical Imaging) has generated encouraging outcomes. In classification, our empirical findings consistently demonstrate that self-supervised training performs better than the conventional supervised approach across Convolutional Neural Network (CNN) and Transformer models.
Moreover, in the comparative analysis of the two self-supervised techniques. CASS demonstrates superiority by exhibiting enhanced performance in nearly all scenarios compared to DINO. Utilizing reduced supervision in segmentation has yielded favorable outcomes. Specifically, applying the Semi-supervised approach, as outlined in [8], has resulted in notable enhancements in performance across all label fractions for the three datasets under consideration.
@@
Based on a comprehensive analysis of our experimental findings, adopting machine-level supervision, such as the S4MI proceses we examined reduces dependence on human supervision and also yields substantial advantages in time and accuracy of medical image analysis. The findings of our study offer significant empirical contributions to medical imaging and limited-supervision techniques, thereby stimulating future investigations in this area. The distribution as open-sourced code of our S4MI processes will support other researchers’ efforts by saving time in labeling and improving image analysis. This, in turn, will facilitate the advancement of healthcare solutions, leading to the eventual enhanced outcomes in patient care. We are optimistic that our study will serve as a catalyst for substantive discourse and cooperative endeavors within the medical imaging community, thereby propelling progress in this pivotal domain.

### v4

@@ 6 Conclusion
The primary focus of our experimental investigations revolved around two fundamental medical imaging tasks: classification and segmentation. Significantly, our findings from employing the S4MI (Self-Supervision and Semi-Supervision for Medical Imaging) framework indicate a promising shift away from traditional, purely transfer-learning-based supervised methodologies. Specifically, in the realm of classification, our empirical evidence consistently shows that self-supervised training within the S4MI framework outperforms conventional supervised approaches across both Convolutional Neural Network (CNN) and Transformer models.
Moreover, in the comparative analysis of the two self-supervised techniques, CASS demonstrates superiority by exhibiting enhanced performance in nearly all scenarios compared to DINO. Utilizing reduced supervision in segmentation has yielded favorable outcomes. Specifically, applying the semi-supervised approach, as outlined in [8], has resulted in notable enhancements in performance across all label fractions for the three datasets under consideration.
@@
Based on a comprehensive analysis of our experimental findings, adopting machine-level supervision through the S4MI framework reduces dependence on human supervision and yields substantial advantages in both time efficiency and accuracy of medical image analysis. The findings of our study make significant empirical contributions to the fields of medical imaging and limited-supervision techniques, thereby stimulating future research in this area. The distribution of our S4MI pipeline as open-sourced code will aid other researchers by saving labeling time and improving image analysis quality. This, in turn, will facilitate the advancement of healthcare solutions, leading to improved patient care outcomes. We are optimistic that our study will act as a catalyst for meaningful dialogue and collaborative endeavors within the medical imaging community, propelling progress in this critical domain.